A movie earned $80 million during its first week of release. Then sales decreased by 60% during each following week. When asked to calculate the amount of money earned during the movie’s 5-week release period, Julian got an answer of about $2 million.
Part A: How can you tell immediately that Julian’s answer is not reasonable?
Part B: What error did Julian likely make?
Part C: How much money did the movie earn during its 5-week release period? Round your answer to the nearest million.

Answer by richwmiller(9144)

a)
We already know that the movie earned $80 million the first week
Five week total of $2 million is lower than one week $80 million.
b)
probably just the last week.
c)
1st week
80 million dollars
2nd week
40% of 80=.4*80=32 million dollars
3rd
40% (.4*80)=.4*.4*80=12.8 million dollars
4th
.4*.4*.4*80=5.12 million dollars
5th
.4*.4*.4*.4*80=2.048 million dollars
total 131.968
total =132 million dollars